# VMTS: Vision-Assisted Teacher-Student Reinforcement Learning for Multi-Terrain Locomotion in Bipedal Robots

### v1

@@ I INTRODUCTION
Figure 2: Training Overview. In the first phase, the network inputs include proprioceptive information, the height map within a one-meter radius around the robot, and accurate physical parameters. Additionally, terrain classification and privileged information are utilized to adjust the weighting of the the mixture of expert, facilitating the training of a reliable teacher policy. In the second phase, the input information consists solely of the current depth map and proprioceptive data. The student policy is trained by aligning the encoded representations of both the student and teacher policies using Barlow Twins, while simultaneously employing supervised learning to guide the student policy to imitate the output of the teacher policy.
@@
  • First, our framework integrates the mixture of experts model with the teacher-student policy, presenting a novel vision-based reinforcement learning framework for legged robots.
  • Second, we introduce an alignment loss between the teacher and student networks, which enhances the performance of the student policy.
  • Finally, we experimentally validate the feasibility and robustness of our approach on a real-world platform.
@@ III-B Environment Setup
We employed NVIDIA Isaac Gym[28] as the simulation environment to train 8192 robots in parallel, with both the teacher and student policies utilizing 8192 robots. The teacher policy converged after approximately 4000 training episodes, requiring around 3 hours on an RTX 3090 GPU. During the training of the student policy, the substantial memory consumption and longer acquisition times for visual information in Isaac Gym significantly increased the training time. Specifically, training the student policy for 4000 episodes, building on the teacher policy, took approximately 120 hours. To address this, we utilized NVIDIA Warp[32] to store terrain mesh data during terrain initialization. When acquiring depth images, we adopted a ray-casting approach based on the robot’s camera pose and physical properties, rather than relying solely on Isaac Gym’s visual data acquisition. This modification, leveraging NVIDIA Warp for depth map retrieval, reduced the training time for the student policy to approximately 16 hours.During the training process, certain training parameters of the robot, such as PD parameters, target base height, foot end height, and others, are configured as shown in Table II.

### v2

@@ I INTRODUCTION
Figure 2: Training Overview. In the first phase, the network inputs include proprioceptive information, the height map within a one-meter radius around the robot, and accurate physical parameters. Additionally, terrain classification and privileged information are utilized to adjust the weighting of the the mixture of expert, facilitating the training of a reliable teacher policy. In the second phase, the input information consists solely of the current depth images and proprioceptive data. The student policy is trained by aligning the encoded representations of both the student and teacher policies using Barlow Twins, while simultaneously employing supervised learning to guide the student policy to imitate the output of the teacher policy.
@@
  • First, our framework integrates a mixture of expert models with teacher-student policy learning, where the teacher strategy adaptively adjusts its policy based on terrain, providing a novel vision-aided reinforcement learning framework for legged robots.
  • Second, compared to previous approaches that directly distill the teacher’s policy to the student, we introduce an implicit alignment and redundancy reduction in the observation feature space between the teacher and student prior to distillation, thereby enhancing the student’s performance.
  • Finally, we propose a smooth foot-trajectory tracking reward function to mitigate the sim-to-real gap, with comprehensive simulation and physical experiments validating the overall efficacy and superiority of our approach.
@@ III-B Environment Setup
We employed NVIDIA Isaac Gym[27] as the simulation environment to train 8192 robots in parallel, with both the teacher and student policies utilizing 8192 robots. The teacher policy converged after approximately 4000 training episodes, requiring around 3 hours on an RTX 3090 GPU. During the training of the student policy, the substantial memory consumption and longer acquisition times for visual information in Isaac Gym significantly increased the training time. Specifically, training the student policy for 4000 episodes, building on the teacher policy, took approximately 120 hours. To address this, we utilized NVIDIA Warp[31] to store terrain mesh data during terrain initialization. When acquiring depth images, we adopted a ray-casting approach based on the robot’s camera pose and physical properties, rather than relying solely on Isaac Gym’s visual data acquisition. This modification, leveraging NVIDIA Warp for depth images retrieval, reduced the training time for the student policy to approximately 16 hours.During the training process, certain training parameters of the robot, such as PD parameters, target base height, foot end height, and others, are configured as shown in Table II.
@@ IV-A Simulation Experiments
Figure 4: The motion timing diagram of a bipedal robot with visual assistance in simulation. The lower section of the diagram illustrates the depth images obtained through NVIDIA Warp and ray casting during the robot’s movement.